OpenStudy (anonymous):

Last year, McBarkie's Burgers contained 200 grams of fat in each burger. This year, the amount of fat has been reduced to 120 grams per burger. What is the percent of decrease in the number of grams of fat per burger to the nearest percent?

OpenStudy (anonymous):

I know that it is not 20%

OpenStudy (anonymous):

percent of decrease = (starting amount - ending amount)/starting amount Last year, the burgers had 200 grams of fat, so that's the "starting amount" This year, they have 120 grams per burger, so that's the "ending amount" Calculate the fraction and express as a percentage...

OpenStudy (anonymous):

What is it, I am stuck on it.

OpenStudy (anonymous):

(start - finish) / start = (200 - 120) / 200 = 80/200 = 40/100

OpenStudy (anonymous):

Divide them now?

OpenStudy (anonymous):

yep, divide 40 by 100... unless you realize what the percentage will be already.

OpenStudy (anonymous):

0.4

OpenStudy (anonymous):

You need 0.4 as a percentage... 0.4 is the same as 4/10, and also the same as 40/100... "forty out of 100"... so it's 40%

OpenStudy (anonymous):

ok. Can you help me on two others

OpenStudy (anonymous):

one other, sorry

OpenStudy (anonymous):

what is it? Post it and try to solve and I can help...

OpenStudy (anonymous):

Change to a decimal: 14 1/2 %.

OpenStudy (anonymous):

that would be 0.0145 right?

OpenStudy (anonymous):

close... if it was 14%, that would be "14 out of 100", or 14/100, or 0.14 So 14.5% would be 14.5 out of 100, or 14.5/100, or 0.145. I think you just got an extra zero in front... 0.0145 is the same as 1 45/100 %

OpenStudy (anonymous):

Thank you.